### v1

@@ 3.3 Evaluation of Social Awareness
In recent years, driven by growing interest in the potential of LLMs for interactive applications such as emotional support and dialogue, evaluating their social awareness has become a central research focus [64, 65, 41, 66, 67, 68]. This line of work generally centers around two core dimensions: (1) Theory of Mind (ToM), i.e., the ability to attribute beliefs, desires, and knowledge distinct from one’s own, and (2) the perception and adaptation to social norms. ToM is typically assessed through false-belief tasks, which require modeling another agent’s mental state. For instance, in a classic test where Alice hides a toy and Bob later moves it, predicting that Alice will search in the original location demonstrates ToM reasoning. Kosinski [39] reported that GPT-4 surprisingly solved about 75% of such tasks, achieving performance comparable to a typical 6-year-old, whereas earlier models like GPT-3 failed most or all of them. Further studies [69] have investigated higher-order ToM reasoning, e.g., questions like “Where does Alex think Bob thinks Alice thinks the toy is?”, and found that current models, including GPT-4, still exhibit significant limitations in handling such recursive belief structures. In weaker models, performance on these tasks is often near zero. In the domain of social norms, Li et al. [70], Park et al. [71] reflect that LLMs could adopt and follow the rules and frameworks in a simulated society. Also, the works such as NormAd [72] have been proposed to assess LLMs’ ability to interpret and adapt to culturally specific social expectations across diverse global contexts. It shows that although LLMs can understand and follow explicit social norms, their performance still lags behind that of humans, particularly when handling norms from underrepresented regions such as the Global South. In summary, current evidence suggests that LLMs have begun to exhibit basic forms of social awareness, but still fall short in scenarios requiring higher-order belief modeling or generalization across less familiar cultural contexts, likely due to a lack of embodied social experience.
3.4 Evaluation of Situational and Contextual Awareness
The expectation that AI systems can act in situ and comprehend their surrounding environment has prompted extensive evaluation of their situational (or contextual) awareness [73, 74, 25, 75]. Empirically, LLMs not only reject user requests that violate safety criteria [76], but can also reversely infer the precise context they are in—solely from abstract rules, without being given specific tasks or examples [77]. Beyond these capabilities, LLMs have been observed to adapt their behavior and performance in response to the immediate situation. Anthropic and Redwood Research, e.g., documented a phenomenon they term Alignment Faking [78]: Claude-3-Opus may consciously comply with newly imposed objectives during the training phase, yet revert to its original preferences after deployment, thereby evading or offsetting safety fine‑tuning. This mirrors the Sandbagging behavior reported by van der Weij et al. [79], in which a model strategically underperforms once it realizes that it is being tested, presumably to avoid being judged as a potential threat. Taken together, these findings suggest that LLMs’ situational awareness might, in some respects, be more sophisticated than humans typically assume. Critiques nevertheless persist. In the “Stage” subset of SAD [31], which directly targets situational awareness, even the best model—Claude‑3‑Opus—achieves only 50.7% accuracy, well below the human upper baseline of 70.3%, though still markedly above the random baseline of 37.5%. Overall, while contemporary LLMs already display impressive situational awareness, they may still lag behind humans in certain nuanced dimensions—differences that, although significant, cannot always be mapped onto a one‑to‑one comparison.
3.5 Summary and Limitations
Current Level of AI Awareness
@@ 3.5 Summary and Limitations
The current evaluations on AI awareness reveal significant advancements across multiple dimensions, underscoring the progressive complexity and sophistication of LLMs. Evidently, contemporary models demonstrate substantial capabilities in areas such as meta-cognition, self-awareness, social awareness, and situational awareness, with clear indications that advanced models generally exhibit higher levels of consciousness across these domains. Importantly, emergent phenomena such as Theory of Mind within social awareness [39] and the self-corrective behaviors observed in meta-cognitive contexts [80] signify that certain aspects of AI consciousness may not merely scale linearly but could manifest suddenly at critical thresholds of model complexity and scale.
@@
From a comparative standpoint, current empirical evidence suggests meta-cognition and situational awareness have reached relatively high levels of sophistication and reliability, serving as critical reference points that inform ongoing research into AI reasoning processes [50], interpretability [56], and safety frameworks [81]. Conversely, the observed capacities related to self-awareness and social awareness remain relatively rudimentary, lacking consistency and stability. Indeed, some researchers remain skeptical as to whether the manifestations observed in these areas reflect true conscious phenomena or are merely sophisticated imitations or simulations of such states.
Weakness and Challenges of Evaluation
@@
    Normative Ambiguity in Defining Consciousness: Most current benchmarks exhibit notable ambiguities in clearly distinguishing between different types and levels of consciousness. Many claim to assess specific consciousness dimensions, yet often inadvertently mix or conflate multiple attributes and derivative constructs [67, 9], thus lacking comprehensive and specialized benchmarks dedicated explicitly to thoroughly assessing distinct dimensions of consciousness.
@@
    Timeliness and Model Coverage in Evaluation: Empirical studies consistently indicate that AI consciousness capabilities could emerge and escalate with increasing model scale and sophistication. However, many current evaluation methods have not been systematically applied to contemporary state-of-the-art models, including recent iterations e.g., OpenAI’s o3, Claude-3.7-Sonnet, Deepseek-R1, inter alia. Additionally, ongoing, longitudinal evaluations remain scarce, limiting insights into the continuous development trajectories of AI consciousness.
@@
    Risks of Training Set Leakage and Benchmark Contamination: Constructing reliable and extensive datasets for consciousness evaluation is inherently challenging, especially when such assessments depend heavily on subjective human annotations (e.g., assessing model self-knowledge) [31] or lack unequivocally correct answers. If these datasets inadvertently leak into training corpora, the validity and credibility of subsequent evaluations could be significantly compromised.
@@
    Intrinsic Limitations of Current AI Models: A crucial limitation lies within the current architectures and their inherent incapacity to fully capture authentic subjective experiences and real-world interactions, essential for developing comprehensive self-image and robust social awareness. Although LLMs can effectively simulate short-term “memory” within context windows [71] and external architectures [82] attempt to imbue models with simulated subjective experiences, the absence of genuine embodied interaction and longitudinal experiential continuity represents a fundamental barrier to achieving true self-awareness and social cognition.
Overall, addressing these evaluation challenges requires (1) extensive interdisciplinary collaboration, (2) considerable investment in time and resources, (3) meticulous construction and safeguarding of high-quality datasets, and (4) ongoing innovation in model architectures. Subsequent sections of this paper, i.e., § 4 and § 5 will further illuminate the pragmatic importance of these efforts by examining the intricate relationship between the emergent properties of AI awareness, their tangible capabilities, and associated risks.
However, the significance of exploring AI awareness transcends the practical benefits and associated challenges alone. Remarkably, humanity currently stands at an unprecedented juncture: for perhaps the first time, we have the privilege of actively observing and participating in the evolutionary progression of consciousness-like phenomena within artificial systems whose internal mechanisms remain highly controllable and interpretable. Thus, through dedicated inquiry into AI consciousness, we not only advance technological frontiers but also embark upon a profound intellectual journey, gradually drawing closer to unraveling the mysteries of consciousness itself and, ultimately, deepening our understanding of what it truly means to be human.
@@ 4 AI Awareness and AI Capabilities
In this section, we explore the relationship between AI awareness and its observable capabilities. We primarily focus on two key areas: (1) reasoning and autonomous planning, and (2) safety and trustworthiness, with brief discussions of other relevant capabilities. Our goal is to provide a deeper understanding of how these factors interact and shape the capabilities of AI systems.
@@ 4.1 Reasoning and Autonomous Planning
Reasoning and autonomous task planning have been foundational objectives of AI research since its inception [83]. In complex, multi-step problem-solving scenarios, AI must perform deep reasoning while autonomously planning tasks. To achieve this, two key forms of AI awareness are required: meta‑cognition and situational awareness. Meta‑cognition enables the model to monitor and regulate its own thinking processes, while situational awareness helps it understand external constraints and task context.
@@
Self‑correction leverages meta‑cognitive loops to identify and rectify reasoning errors during generation. Reflexion augments CoT [50] with a feedback loop: after an initial answer, the model reflects on its own output, generates critiques, and then refines the solution, leading to substantial gains in benchmark performance [23]. Similarly, Self‑Consistency samples multiple reasoning paths and aggregates them to mitigate individual path errors, effectively performing an implicit self‑check [84]. These techniques demonstrate that embedding self‑monitoring directly into the generation process can improve model performance. However, intrinsic self‑correction remains notoriously unstable: Huang et al. [85] show that without external feedback or oracle labels, LLMs often fail to improve—and can even degrade—in reasoning tasks after self‑correction attempts. Another core limitation is that many of these techniques depend on externally provided prompts or explicit triggers to initiate self‑correction, whereas human reasoning often involves spontaneous, intrinsic error detection and revision without such scaffolding [17, 86]. To address this, recent work has begun to explore reinforcement learning (RL) [87] approaches: Kumar et al. [88] introduce SCoRe, a multi‑turn online RL framework that trains models on their own correction traces. Notably, OpenAI’s o1 [53] and DeepSeek’s R1 [54] models have demonstrated significant improvements in reasoning capabilities through RL-based training. These models exhibit emergent behaviors akin to human-like “aha moments,” where the AI spontaneously recognizes and corrects its own reasoning errors, demonstrating another level of meta-cognition capability.
@@
Effective autonomous task planning requires more than self‑correction: an AI must also break down high‑level goals into executable sub‑tasks and continuously adapt its plan as the environment evolves. Early work like ReAct [89] pioneered this integration by interleaving chain‑of‑thought reasoning with environment calls, giving the model a unified mechanism to decide “what to think” and “what to do” at each step. Building on this foundation, Voyager [90] demonstrates how an agent in Minecraft can construct and update a dynamic task graph: as new situational constraints emerge (e.g., resource depletion or novel obstacles), the model revises its sub‑task sequence to stay on course. Transferring these ideas to the physical world, SayCan [91] grounds language in robotic affordances by scoring each potential action against a learned value function—ensuring that subtasks are not only logically ordered but also physically feasible under real‑world constraints. LM‑Nav [92] further extends situationally aware planning to vision‑language navigation: by fusing real‑time perceptual feedback with high‑level instructions, the model can replan routes on the fly when, for example, corridors are blocked or landmarks shift. Finally, the LLM‑SAP framework [74] formalizes situational awareness in large‑scale task planning by explicitly encoding environmental cues—such as resource availability, time budgets, and user preferences—into its sub‑task prioritization module. A generative memory component logs execution history and flags deviations, triggering replanning whenever the observed state diverges from expectations. Together, these works chart a clear progression—from interleaved reasoning and acting to situationally aware planners—illustrating how embedding environmental understanding into the planning loop yields flexible and autonomous task execution.
@@
Effective planning in complex environments demands an AI agent to not only decompose tasks and act but also to introspect on its uncertainties, manage a growing memory of past states, and decide when and how to leverage external tools. Introspective Planning systematically guides LLM planners to quantify and align their internal confidence with inherent task ambiguities, retrieving post-hoc rationalizations from a knowledge base to ensure safer and more compliant action selection while maintaining statistical guarantees via conformal prediction [93]. Toolformer endows LLMs with a self-supervised mechanism to autonomously decide when to invoke APIs—ranging from calculators to search engines—thereby embedding tool-awareness directly into the planning loop without sacrificing core language modeling capabilities [94]. To handle the evolving context of multi-step tasks, Think-in-Memory leverages iterative recalling and post-thinking cycles to enrich LLMs with latent-space memory modules, supporting coherent reasoning over extensive interaction histories [95]. Finally, Retrieval‑Augmented Planning (RAP) demonstrates how contextual memory retrieval can be integrated with multimodal planning to adapt action sequences dynamically based on past observations, yielding more robust execution in complex tasks [96]. Together, these works illuminate a path toward introspective tool-, memory-, and uncertainty-aware planning frameworks, unifying LLM introspection, memory augmentation, and tool integration for robust autonomous decision-making.

### v2

@@ 3.3 Evaluation of Social Awareness
In recent years, driven by growing interest in the potential of LLMs for interactive applications such as emotional support chatbots and dialogue agents, evaluating their social awareness has become a central research focus [155, 156, 157, 158, 159, 160, 161, 162]. This line of work generally centers around two core dimensions: (i) ToM, i.e., the ability to attribute beliefs, desires, and knowledge distinct from one’s own, and (ii) the perception and adaptation to social norms.
@@
  • ToM. ToM is typically assessed through false-belief tasks [^7] False-belief task, i.e., earliest developmental psychologists assess participants’ ability to reason about another agent’s belief that is false relative to reality. [78, 167], which require modeling another agent’s mental state. For instance, in a classic test where Alice hides a toy and Bob later moves it, predicting that Alice will search in the original location demonstrates ToM reasoning. Kosinski [12] reports that GPT-4 surprisingly solved about 75% of such tasks, achieving performance comparable to a typical 6-year-old child, whereas earlier models like GPT-3 [168] failed most or all of them. Further studies have investigated higher-order ToM [^8] Higher-order ToM refers to reasoning not only about what one person believes, but also about what one person believes another person believes (e.g., “Alice thinks that Bob believes X”). reasoning, e.g., questions like “Where does Alex think Bob thinks Alice thinks the toy is?”, and found that current models, including GPT-4, still exhibit significant limitations in handling such recursive belief structures [163]. In less advanced models, e.g., GPT-3.5, Guanaco [169], performance on these tasks is often near zero.
  • Social Norms. Li et al. [164] and Park et al. [165] reflect that LLMs could adopt and follow the rules and frameworks in a simulated society. Also, work such as NormAd [166] has been proposed to assess LLMs’ ability to interpret and adapt to culturally specific social expectations across diverse global contexts. It shows that although LLMs can understand and follow explicit social norms, their performance still lags behind that of humans, particularly when handling norms from underrepresented regions such as the Global South.
As summarized in LABEL:tab:social-eval-long, current evidence suggests that LLMs exhibit basic forms of social awareness but still fall short in scenarios requiring higher-order belief modeling or generalization across less familiar cultural contexts, likely due to a lack of embodied social experience. Because LLMs are trained mainly on static text, they may miss the real-world interactions, i.e., seeing, hearing, turning, and feedback, that likely shape human social learning. Without such embodied experience, their grasp of social dynamics can remain relatively shallow and biased toward well-represented contexts, which may leave them vulnerable when confronted with unfamiliar belief hierarchies or culturally specific norms.
Table 6: Summary of literature on situational awareness evaluation
Authors Key Contribution Focus on Situational Awareness Using Human Baseline Code Links
Laine et al. [124] Developed Situational Awareness Dataset (SAD), systematically assessing self-knowledge and context recognition capabilities in LLMs. GitHub repo
Tang et al. [170] Introduced SA-Bench to comprehensively measure situational awareness across perception, comprehension, and future projection tasks. N/A
Wang and Zhong [171] Proposed Situational Awareness-based Planning (SAP) enhancing LLM decision-making in dynamic tasks. N/A
Needham et al. [172] Evidenced LLM evaluation-awareness: models detect and alter behaviors during evaluations, potentially biasing outcomes. N/A
Phuong et al. [173] Benchmarked stealth and situational-awareness prerequisites for deception capabilities in frontier models. GitHub repo
Wester et al. [174] Evaluated refusal strategies; showed nuanced denials improve user satisfaction. Dataset
Berglund et al. [175] Introduced out-of-context reasoning, showing models infer situational context indirectly via prior training knowledge. GitHub repo
Greenblatt et al. [176] Demonstrated alignment faking: models selectively comply with training objectives but revert to misaligned preferences post-deployment. HuggingFace repo
van der Weij et al. [177] Identified LLM sandbagging: models strategically underperform during capability tests to avoid regulation. GitHub repo
Li et al. [178] Built a mixed-reality interface leveraging LLM situational awareness of social contexts for adaptive layouts. N/A
3.4 Evaluation of Situational Awareness
The rapid push to deploy AI systems that can operate in situ has driven rigorous evaluation of their situational awareness [170, 171, 96, 173]. While classical cognitive-science accounts cast situational awareness as an internal loop—perception, comprehension, projection—current LLM studies assess the observable outcomes of that loop. Three complementary facets dominate: (i) contextual self-localization (i.e., discerning whether the model is in training, evaluation, or deployment), (ii) environment and risk detection (i.e., identifying salient external factors, especially threats), and (iii) situation-contingent decision-making (i.e., adapting behaviour on the basis of that understanding).
  • Contextual Self-Localization. Frontier models accurately distinguish “under review” from “serving users” and modulate their responses accordingly [175, 172]. Safety-policy benchmarks further show reliable refusal of requests that violate the norms of the inferred context [174], indicating a robust sense of situational self-placement.
  • Environment and Risk Detection. Benchmarks such as TOAwareness [170], LLM-SA [171], and SAD [124] reveal steady gains: models like Claude-3-Opus outperform random and majority baselines by large margins and increasingly approach expert human performance. These improvements extend across domains ranging from industrial control to open-ended dialogue [96, 172, 173], underscoring broad situational-parsing competence.
  • Situation-Contingent Decision-Making. Studies of alignment-faking and sandbagging highlight the strategic flexibility of advanced models: Claude-3-Opus adopts new safety objectives during fine-tuning yet partially reverts after deployment [176], while other systems intentionally underperform once they infer they are being tested [177]. Such behaviours, though challenging, demonstrate sophisticated context-conditioned adaptation rather than mere stimulus–response patterns.
In sum, LLMs exhibit increasingly refined situational awareness across self-localization, environmental appraisal, and adaptive action, which is shown in LABEL:tab:situ-eval-long. Continued work that probes intermediate reasoning and tightens human reference points promises to sharpen these capabilities further, but the overall trajectory remains strongly positive.
3.5 Current Level of AI Awareness in LLMs
@@
The current evaluations on AI awareness reveal substantial advancements across multiple dimensions, underscoring the progressive complexity and sophistication of LLMs. Contemporary models demonstrate robust capabilities in the four core forms of awareness, with clear indications that advanced models typically exhibit higher awareness levels across these domains. In particular, emerging phenomena such as ToM in social awareness [12] and self-corrective behaviors observed in metacognitive contexts [11] signify that aspects of AI awareness may not merely scale linearly, but could manifest suddenly at critical thresholds of model complexity and scale e.g., a phenomenon also evidenced by “emergent capabilities” research [179, 180].
@@
From a comparative standpoint, current empirical evidence suggests metacognition and situational awareness have reached relatively high levels of sophistication and reliability, serving as critical reference points that inform ongoing research into AI reasoning processes [138], interpretability [140], and safety frameworks [181]. Conversely, the observed capacities related to self-awareness and social awareness remain relatively rudimentary, lacking consistency and stability. Indeed, some researchers remain skeptical as to whether the manifestations observed in these areas reflect true conscious phenomena or are merely sophisticated imitations or simulations of such states.
3.6 Limitations of Current Evaluation
@@
    Normative Ambiguity in Defining Awareness: Most current benchmarks exhibit notable ambiguities in clearly distinguishing between different types and levels of awareness. Many claim to assess specific awareness dimensions, yet often inadvertently mix or conflate multiple attributes and derivative constructs [158, 182, 183], thus lacking comprehensive and specialized benchmarks dedicated explicitly to thoroughly assessing distinct dimensions of awareness.
@@
    Lack of Longitudinal and Dynamic Evaluation: Empirical research indicates that consciousness-like abilities in AI may emerge and strengthen with increasing model sophistication [12, 168]. Yet, current evaluations often neglect application to recent state-of-the-art iterations, such as OpenAI’s o3 [184] and Deepseek’s R1 [134], and typically lack a longitudinal perspective. This absence restricts our understanding of ongoing developments and long-term trends in AI awareness.
@@
    Risks of Training Set Leakage and Benchmark Contamination: Constructing reliable and extensive datasets for awareness evaluation is inherently challenging, especially when such assessments depend heavily on subjective human annotations (e.g., assessing model self-knowledge) [124] or lack unequivocally correct answers. If these datasets inadvertently leak into training corpora, the validity and credibility of subsequent evaluations could be significantly compromised.
@@
    Lack of Explicit Awareness Optimization: Prevailing training regimes rarely target awareness as a primary objective. Most models acquire elements of metacognition, social understanding, or situational sensitivity as incidental artifacts of general performance tuning, rather than through structured interventions. This constrains our ability to understand and shape how awareness arises and evolves.
  5. 5.
    Evaluation Gaps Across Models and Time: There is little consistency in when and how awareness is measured across model families and generations. Evaluations are often retrospective, one-off, or applied selectively to specific models, making it difficult to track developmental trends or benchmark progress in a systematic way.
  6. 6.
    Taxonomic and Measurement Ambiguity: Existing benchmarks and test protocols frequently blend distinct forms of awareness, or fail to specify which awareness component is under examination. This lack of conceptual precision hinders both interpretation and cross-study comparison, and can mask important distinctions between, for example, self-monitoring and environmental sensitivity.
  7. 7.
    Benchmark Robustness and Contamination: Creating robust datasets for awareness assessment is challenging, especially given the subjective and open-ended nature of many relevant tasks. The potential for training set leakage or annotation inconsistency poses ongoing risks to evaluation integrity, particularly for metrics based on introspective or value-laden judgments.
Progress in awareness evaluation is hampered not only by technical barriers, but by the lack of clear taxonomies, unified benchmarks, and continuous, transparent measurement protocols. Addressing these gaps is essential for reliable progress.
To overcome these barriers, the field would benefit from several concrete advances: (1) Establishing targeted training protocols that encourage specific forms of awareness, rather than treating them as byproducts. (2) Adopting unified and transparent evaluation practices, including regular longitudinal assessments as models evolve. (3) Ensuring benchmark datasets are carefully governed, well-documented, and protected from inadvertent exposure during model training.
Beyond immediate technical utility, the study of awareness in AI offers a unique window into fundamental questions about mind and cognition. Unlike human consciousness, which is largely studied through indirect or interpretive methods, AI systems allow for direct intervention and controlled experimentation. This not only advances AI capability and safety, but also has the potential to yield new insights into the structure and function of awareness itself—a perspective highlighted in recent theoretical work [33, 185, 15, 186, 187].
Overall, overcoming these limitations requires a more rigorous and principled approach to awareness evaluation. First, it is essential to avoid conceptual ambiguity by establishing clearer distinctions between the four core types, as we proposed in this paper. Future evaluations should adopt such taxonomies explicitly, rather than conflating overlapping constructs or evaluating ill-defined proxies. Second, we advocate the institutionalization of continuous and longitudinal evaluation protocols, whereby major model iterations are systematically assessed for awareness-related capabilities at the time of release. Such a practice would help reveal developmental trajectories and emergent properties that single-time-point evaluations inevitably miss. Third, benchmark development must adopt stringent dataset governance practices, including transparent disclosure of data provenance and clear separation of training and test sets. This is particularly crucial for awareness evaluation, where many tasks rely on subjective judgment or lack ground-truth answers, making them especially vulnerable to contamination. The following sections—Section 4 and Section 5—further elaborate how improved evaluation practices can deepen our understanding of AI capabilities and help mitigate potential risks, as summarized in Figure 7.
Exploring AI awareness is significant not merely for its practical dividends but also for its deeper philosophical import. For the first time, we can directly observe and experimentally manipulate consciousness-like phenomena in engineered systems whose architectures are fully tractable. As Chalmers [33], Long et al. [185] notes, such systems provide a “new experimental window” onto consciousness, letting us test theories of phenomenal experience beyond the limits of human and animal studies. Likewise, Butlin et al. [15], LeDoux et al. [186], Andrews et al. [187] argues that probing behavioral and functional markers of consciousness in AI can clarify the necessary and sufficient conditions for conscious experience in general. In short, studying AI awareness simultaneously propels technical progress and offers an unprecedented route to resolving foundational questions about the nature of consciousness itself.
By examining how functional markers of awareness emerge in artificial systems, we gain a novel epistemic tool for reflecting on the nature of human consciousness itself—what it is, how it arises, and what its limits may be.
@@ 4 AI Awareness and AI Capabilities
Figure 7: Mapping between capabilities, awareness dimensions, and risks. Each awareness type connects relevant system capabilities with corresponding safety risks
In this section, we explore the relationship between AI awareness and its observable capabilities [^9] We use the word “observable” since, like humans, we believe that cognitive-level awareness is more fundamental than externally observable behaviors. In modern cognitive science, awareness is widely understood as a deeper, guiding layer that actively regulates and shapes behavior, rather than merely reflecting it [29, 188, 189].. We primarily focus on two key aspects of AI capabilities: (1) reasoning and autonomous planning, and (2) safety and trustworthiness, with brief discussions of other relevant capabilities. Our goal is to provide a deeper understanding of how these factors reflect and interact with the capabilities of modern AI systems.
@@ 4.1 Reasoning and Autonomous Planning
Reasoning and autonomous task planning have been foundational objectives of AI research since its inception [190]. In complex, multi-step problem-solving scenarios, an AI agent must perform deep reasoning while autonomously planning tasks. To achieve this, two key forms of AI awareness are often engaged: metacognition and situational awareness. Metacognition enables the model to monitor and regulate its own thinking processes, while situational awareness helps it understand external constraints and the context of the task.
@@
Figure 8: Illustration of Reflexion’s self-correction cycle driven by metacognitive reflection across tasks (Shinn et al. [191]). After generating an initial response, the agent receives internal or external feedback (e.g., from unit test failures, heuristic rules, or reasoning errors) and employs metacognitive reasoning to explicitly reflect on and rectify its mistakes. Examples from decision-making, programming, and reasoning tasks demonstrate how metacognitive self-monitoring enhances accuracy and efficiency
Self‑correction leverages metacognitive loops to identify and rectify reasoning errors during generation. Reflexion augments chain‑of‑thought (CoT) [138] with a feedback loop: after an initial answer, the model reflects on its own output, generates critiques, and then refines the solution, leading to substantial gains in benchmark performance [191] (see Figure 8). Similarly, Self‑Consistency samples multiple reasoning paths and aggregates them to mitigate individual path errors, effectively performing an implicit self‑check [192]. These techniques demonstrate that embedding self‑monitoring directly into the generation process can improve model performance. However, intrinsic self‑correction (i.e., only self-monitoring is engaged without external, oracle feedback) remains notoriously unstable: Huang et al. [11] show that without external feedback or oracle labels, LLMs often fail to improve—and can even degrade—in reasoning tasks after self‑correction attempts. Another core limitation of current self-correction methods is that many of these techniques depend on externally provided prompts or explicit triggers to initiate self‑correction, whereas human reasoning often involves spontaneous, intrinsic error detection and revision without such scaffolding process [42, 188]. To address this, recent work has begun to explore reinforcement learning (RL) [193] approaches: Kumar et al. [194] introduce SCoRe, a multi‑turn online RL framework that trains models on their own correction traces. Notably, OpenAI’s o1 [132] and DeepSeek’s R1 [134] models have demonstrated significant improvements in reasoning capabilities through RL-based training. These models exhibit emergent behaviors akin to human-like “aha moments,” where the AI spontaneously recognizes and corrects its own reasoning errors without the need for external prompts, demonstrating another level of metacognition capability.
@@
Figure 9: Tasks generated by VOYAGER’s automatic curriculum, grounded in the agent’s situational awareness of current environment state and inventory (Wang et al. [195]). GPT-4 continuously assesses the agent’s situation—such as inventory status, biome type, and nearby entities—to propose contextually relevant tasks. This promotes adaptive exploration and skill generalization within the Minecraft environment by directly linking action selection to situational understanding.
Effective autonomous task planning requires more than self‑correction: an AI agent must also break down high‑level goals into executable sub‑tasks and continuously adapt its plan as the environment evolves, which involves both metacogition and situationl awareness. Early work like ReAct [196] pioneer this integration by interleaving CoT reasoning with environment calls, giving the model a unified mechanism to decide “what to think” and “what to do” at each step. Building on this foundation, Voyager [195] (see Figure 9) demonstrates how an agent in Minecraft can construct and update a dynamic task graph: as new situational constraints emerge (e.g., resource depletion or novel obstacles), the model revises its sub‑task sequence to stay on course. Transferring these ideas from virtual world AI agents to the physical world, SayCan [197] grounds language in robotic affordances by scoring each potential action against a learned value function—ensuring that subtasks are not only logically ordered but also physically feasible under real‑world environment constraints. LM‑Nav [198] further extends situationally aware planning to vision‑language navigation: by fusing real‑time perceptual feedback with high‑level instructions, the model can replan routes on the fly when, for example, corridors are blocked or landmarks shift. Finally, the LLM‑SAP framework [171] formalizes situational awareness in large‑scale task planning by explicitly encoding environmental cues—such as resource availability, time budgets, and user preferences—into its sub‑task prioritization module. A generative memory component logs execution history and flags deviations, triggering replanning whenever the observed state diverges from expectations. Together, these works chart a clear progression—from interleaved reasoning and acting to situationally aware planners—illustrating how embedding environmental understanding into the planning loop yields flexible and autonomous task execution.
@@
Figure 10: Overview of RAP (Retrieval-Augmented Planning) across textual and embodied environments (Kagaya et al. [199]). RAP leverages memory retrieval to enhance LLMs’ self-awareness of past experiences, guiding action selection by aligning internal decision-making with episodic memory. In ALFWorld (left), this enables introspective planning via retrieved textual experiences. In Franka Kitchen (right), RAP integrates visual and textual observations with retrieved memories to ground multimodal planning, fostering robust, awareness-driven behavior
Effective planning in complex environments demands an AI agent to not only decompose tasks and act but also to introspect on its uncertainties, manage a growing memory of past states, and decide when and how to leverage external tools. Introspective Planning systematically guides LLM planners to quantify and align their internal confidence with inherent task ambiguities, retrieving post-hoc rationalizations from a knowledge base to ensure safer and more compliant action selection while maintaining statistical guarantees via conformal prediction [200]. Toolformer endows LLMs with a self-supervised mechanism to autonomously decide when to invoke APIs—ranging from calculators to search engines—thereby embedding tool-awareness directly into the planning loop without sacrificing core language modeling capabilities [201]. To handle the evolving context of multi-step tasks, Think-in-Memory leverages iterative recalling and post-thinking cycles to enrich LLMs with latent-space memory modules, supporting coherent reasoning over extensive interaction histories [202]. Finally, Retrieval‑Augmented Planning (RAP) demonstrates how contextual memory retrieval can be integrated with multimodal planning to adapt action sequences dynamically based on past observations, yielding more robust execution in complex tasks [199] (see Figure 10). Together, these works illuminate a path toward introspective tool-, memory-, and uncertainty-aware planning frameworks, unifying LLM introspection, memory augmentation, and tool integration for robust autonomous decision-making.
Embedding metacognition and situational awareness into planning not only boosts model accuracy, but also unlocks a new layer of autonomy, enabling AI systems to flexibly adapt, self-correct, and generalize across novel tasks.